OpenStudy (rebeccapink6398):

WILL MEDAL!!!! Which ratio correctly compares 4 feet to 36 inches?

OpenStudy (anonymous):

1 feet = 12 inches, 4 feet= 48 inches 4feet/36 inches= 48/36=4/3=1.33

OpenStudy (rebeccapink6398):

A.1/10 B.1/9 C.3/4 D.4/3

OpenStudy (rebeccapink6398):

@Bella5 could you help? :D @appleboy2 went off online

OpenStudy (anonymous):

D 4/3 :) Convert both to the same measure, and you have your answer. 4 feet is 48 inches, so it could be 48:36. But 36 inches is also 3 ft, so a simpler ratio would be 4:3. 4:3 is the same as 4/3.

OpenStudy (rebeccapink6398):

THX!!!